OpenStudy (anonymous):

A playground is being designed where children can interact with their friends in certain combinations. If there is 1 child, there can be 0 interactions. If there are 2 children, there can be 3 interactions. If there are 3 children, there can be 12 interactions. If there are 4 children, there can be 39 interactions. Which recursive equation represents the pattern?

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

an = an – 1 + 3n – 2 an = an – 1 + 2n – 3 an = an – 1 + (3n – 2) an = an – 1 + (2n – 3)